Shift report 20030306 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Mar 06 2003 - 00:01:42 EST

  • Next message: Apache: "Shift report 20030306 00:00-08:00"
    ------------------------------------------------------------
    Date: Thu Mar  6 00:01:42 2003, Shift: 16:00-24:00
    Supervisor: ykl, with: rw
    Log book#14 pages 47 - 47
    
    Shift Summary:
    ------------------------------------------------------------
    1. After the access to IR area MCR continued to inject with 6 bunch ramp
     without success.
    
    2. No achievement except pedestal run (8209), TPC pedestal run (8210), and
    DVM calibration run (8211).
    
    3. T3 is touchy. We had no chance to turn it on but FV suggested that if
    any anode trips leave that section (A,C,D) turned off.
    
    Plan: MRS @90 deg,A1050,for "lots of statistics (FV)".
          FS @4 deg, 1/4B for 200K.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Mar 06 2003 - 00:02:27 EST